a parallelogram and a rhombus are equal in area.the diagonals of the rhombus are 110m and 44 m.if one of the sides of parallelogram is 88m,find its corresponding altitude
Answer by dkppathak(439)

area of rhombus =1/2d1xd2
=1/2x110x44
=2420 sq meter
area of parallelogram is equal to rhombus
area of parallelogram is equal=2420 sq meter
area of 1/2 of parallelogram=side x corresponding altitude
88xh=2420
h=2420/88=27.5 meters
answern 27.5 meters